OpenStudy (pulsified333):

Use linear approximation to approximate √36.4 as follows. Let f(x)=√x. The equation of the tangent line to f(x) at x=36 can be written in the form y=mx+b. Compute m and b. m= b= Using this find the approximation for √36.4. Answer:

OpenStudy (anonymous):

hi

OpenStudy (pulsified333):

hi

OpenStudy (anonymous):

f(x)=f(a) +f´(a) (x-a) f(a) = sqrt(36)=6 f´(a) =1/2sqrt36 = 1/12 so f(36.4) = 6 +1/12 (36.4-36) = 6+0.0333 =6.0333 The real value is 6.033241252 so why would wed use 9 significant figures

OpenStudy (pulsified333):

@Ghelper81 im a little confused

OpenStudy (loser66):

why?